A child of mass m is released from rest at the top of a water slide, at height h = 850.0 cm above the bottom of the slide. Assuming that the slide is frictionless because of the water on it, the child’s speed at the bottom of the slide is…..
1
Expert's answer
2020-09-02T12:14:21-0400

As per the given question,

Mass of the children=m

Height of the water slide"(h)=850cm =8.5m"

Let the gravitational acceleration is "(g)=9.8m\/sec^2"

The speed of the child, at the bottom of the slide =v

Initial velocity of the child=u,

"v^2=u^2+2gh"

Now, substituting the values,

"\Rightarrow v^2=0+2\times 9.8\times 8.5m"

"\Rightarrow v=\sqrt{166.6} m/sec"

"\Rightarrow v=12.90 m/sec"